OpenStudy (anonymous):

Multiply and simplify, if possible. (sq. rt. 2 +2)(sq. rt. 2-5) Please walk me thru:)

OpenStudy (anonymous):

(sq. rt. 2 +2)(sq. rt. 2-5) Expand it first (sqrt2)(sqrt2)-5(sqrt2)+2sqrt2 -2(5) =2-3sqrt2-10 <-- simplify it =-8-3sqrt2 <-- group terms

OpenStudy (anonymous):

The FOIL property (First Outer Inner Last) (a+b)(c+d) = ac + ad + bc + bd. Notice that a and c are the FIRST in their respective brackets, a and d are the outer guys, b and c are the INNER guys, and b and d are the last.

OpenStudy (anonymous):

Thanks..CoCo Tsoi and mareoraft!!!

OpenStudy (anonymous):

you are welcome :D